# LLM-Enhanced Multi-Agent Reinforcement Learning with Expert Workflow for Real-Time P2P Energy Trading

### v2

@@ III-B1 Preparation
The expert strategy constraints are embedded within a multi-agent actor-critic framework to solve the constrained optimization problem. Based on the Lagrangian dual theory, this problem is first transformed into its Lagrangian dual form. For each prosumer agent $i$, the multi-agent formulation is can be expressed as:
@@
where $\mathcal{B}$ is the experience replay buffer; $\phi$ is the actor network parameters and the output sampling from the Gaussian distribution $a_{i,t}\sim\mathcal{N}(\mu_{\phi_{i}},\sigma_{\phi_{i}}^{2})$; $\epsilon$ is the policy deviation; $\theta$ is the Q-function network parameters and z = 1,2; $\psi$ is the V-function network parameters. The policy network outputs a distribution over actions, which inherently facilitates exploration during training by enabling diverse action sampling under identical states.
@@
The resulting Wasserstein-2 metric between the expert strategy and the actor network is:
@@ III-B2 Multi-Head Differential Attention
To enhance the modeling capability of the centralized critic in capturing inter-agent interactions, a multi-head differential attention mechanism is integrated into the critic network. This design aims to improve the accuracy of global value estimation. For each agent $i$, the input is first transformed into an embedding $e_{i}$ and then concatenated into the resulting matrix $E$ via an input embedding module, which is implemented as a multi-layer perceptron (MLP). Matrix $E$ is then split into $h$ attention heads and linearly projected into the query, key, and value spaces for each head, as follows:
@@
To further enhance the representational capacity and mitigate issues such as gradient vanishing and feature degradation, a residual network structure is employed following the attention module. This design preserves input information and facilitates efficient training. The intermediate output $X$ is then concatenated with the original input $E$ and the combined vector is processed through another MLP to obtain the final output value. The architecture are shown in the figure 9.
@@ III-B3 Learning the Critics
In scenarios where agents share a global reward, a major challenge is to reduce the variance of policy gradient estimates in interactive multi-agent environments. To address this issue, a double Q-function network and a target V-function network are employed. The minimum selection operation in the double Q-function network effectively mitigates overestimation bias. The incorporation of a V-function network contributes to a significant reduction in estimation variance [32].
For V-function network $V_{\psi_{i}}$, which the loss function is calculated approximates the given the current state and Q-function values:
$\mathcal{L}_{V}(\psi_{i})=\mathbb{E}_{s_{t}\sim\mathcal{B},{a}_{t}\sim\pi_{\phi_{i}}(\cdot|{{o}_{i,t}})}\Bigl[(V_{\psi_{i}}({s}_{t})-\underset{z\in{1,2}}{\text{min}}Q_{\theta_{i,z}}(s_{t},{a}_{t}))^{2}\Bigr].$ (39)
For the Q-function network $Q_{\theta_{i}}$, the training target $y_{t}$ is computed using the a delay updated target network $V_{{\overline{\psi}}_{i}}$. The loss function is defined as follows:
$\begin{split}\mathcal{L}_{Q}(\theta_{i})=&\mathbb{E}_{(s_{t},\mathbf{a}_{t},r_{t},s_{t+1})\sim\mathcal{B}}\biggl[\frac{1}{2}\left(Q_{\theta_{i}}(s_{t},a_{t})\right.\left.-y_{t}\right)^{2}\biggr]\\ &y_{t}=r_{i,t}+\gamma V_{\bar{\psi}_{i}}(s_{t+1})\ ,\end{split}$ (40)
where the target V-function network ${\overline{\psi}}_{i}$}is updated via Polyak averaging rather than direct copying to enhance stability:
$\bar{\psi}_{i}\leftarrow\tau\psi_{i}+(1-\tau)\bar{\psi}_{i}\ ,$ (41)
@@ III-B4 Learning the Actors
In the initial phase of training, a low policy deviation is employed to guide the agent’s learning; during the middle and later stages, a larger policy deviation is introduced to sustain the agent’s exploration. This framework ultimately enables efficient and stable policy optimization.
@@ IV NUMERICAL STUDY
To validate the effectiveness of the proposed framework, a numerical study is carried out on a modified IEEE 141‑bus distribution networks, the voltage level is 12.47kV. As shown in Fig.4, 20 prosumers are selected to participate in local P2P energy trading, and the characteristics of five prosumer types are summarized in TableI. The renewable generation outputs and load demand curves of the prosumers are extracted from a real‑world dataset published by the Belgian Transmission System Operator Elia[34].
Node Devices Portfolio Prosumer Scenario CDG WT PV ESS CL 48,78,102,127, Commercial 59,109,130,140 Rural 67,95,133,136 Industrial 62,86,106,138 Residential 74,100,116,134 Energy Hub
TABLE I: Configurations of personalized prosumers
@@ IV-A Comparison Baselines
    MAGAIL[38]: A multi-agent adaptation of GAIL that uses an adversarial network to imitate expert trajectories provided by the LLM.
@@
    Our Proposed: Introduces a Lagrange multiplier during training to progressively constrain agent behaviors toward the LLM expert’s actions.
@@ IV-B Implementation Details
In terms of neural network architecture design, aside from the variants incorporating attention mechanisms, all baseline algorithms share a unified network architecture under identical hyperparameter settings. The detailed hyperparameter settings are provided in TABLE II. Training was conducted for 5000 episodes with the same random seed initialization and update frequency. All experiments were implemented in Python 3.11.10 under the PyTorch 2.7, with parameters updated via the Adam optimizer. Computations were performed on a platform equipped with an NVIDIA RTX 5070Ti GPU, an AMD Ryzen ThreadRipper 3970X CPU, and 64GB of RAM.
TABLE II: HYPERPARAMETERS USED IN THE MARL
@@
Using year-round data, we designate the first day of each month as the validation set, the last week of each month as the test set, and the remaining days for training. A dynamic validation strategy is employed during training. After each training step, the current policy is evaluated using the validation set. The average cumulative discounted reward across all validation sets is recorded as the performance metric. To ensure statistical robustness, each algorithm is executed independently five times, and both the mean award and standard deviation of the resulting rewards are recorded.
In the present study, the LLM workflow is implemented through online API interfaces for experimental validation. Nevertheless, the proposed architecture is fully modular, allowing seamless substitution with locally hosted or SFT LLM instances to ensure data privacy and reduce latency. The LLM workflow is activated only during the training or system reconfiguration stages (e.g., when new devices are added or topology changes occur). During the training phase, each prosumer executes the LLM pre-generated optimization model code in parallel, while the DSO merely aggregates the objective functions and constraints from all prosumers and utilizes a commercial solver to generate secure operational strategies for each prosumer. In real-time operation, the lightweight MARL agents autonomously perform decision-making without further LLM guidance or global DSO validation. Consequently, the proposed framework achieves high computational efficiency and scalability, ensuring that decision-making comfortably satisfies the 15-minute market interval requirement even when extended to large-scale prosumer networks.
@@ IV-C1 Analysis of the LLM Expert Strategy
In this evaluation, the human expert baseline refers to the optimal strategy obtained by solving a deterministic convex optimization model—formulated in accordance with the problem setting described in Section II—using the Gurobi commercial solver. This model fully incorporates all physical constraints of the prosumers and minimizes the total operational cost under perfect forecasts of renewable generation and load demand. The resulting solver-based optimal policy serves as the ground-truth reference for assessing the accuracy of LLM-generated strategies.
@@
TABLE III: Performance comparison of different LLMs in workflow
@@
TABLE III provides a detailed comparison of the four metrics proposed in this study, where LLMs above the dashed line deactivate advanced reasoning capabilities, while those below activate it. The latency of a large model depends on network speed when API calls are used, but it is effectively eliminated when the model is deployed locally. Specifically, LLMs without advanced reasoning capabilities are generally more cost-efficient; among these, Google Gemini-2.5-Flash shows the highest average cost at 2.43 CNY (Derived from the official billing statements of different LLMs after running the workflow once, and converted into CNY based on the prevailing exchange rate), mainly because it generates a larger number of output tokens per query, leading to higher overall billing despite its relatively low per-token price. In contrast, among reasoning-enabled models, Claude 4-Opus is the most expensive, with an average price of 6.27 CNY, primarily due to its higher per-token pricing structure reflecting the premium positioning of its advanced reasoning capabilities. Open-source LLMs such as DeepSeek and Qwen can further eliminate API-related expenses entirely when deployed locally.
@@
Figure 5: Comparative experiments on the framework proposed in this paper
@@ IV-C2 Performance Analysis of MARL Algorithm
The results in Fig. 6 demonstrate the performance of the proposed algorithm during training on the validation set. As shown, ’Our proposed-MH’ achieves faster convergence to a low operational cost with minimal fluctuations, demonstrating enhanced reward stability for guiding agents when LLM outputs expert actions. Furthermore, the curve of average voltage violation rate for ’Our proposed-MH’ rapidly declines and maintains a low level, highlighting its strength in constraint satisfaction and ability to maintain secure grid operations.
TABLE IV: Comparison of different algorithms on test set
@@
After completing model training, we applied the proposed algorithm to an test set to evaluate the practical applicability. As shown in TABLE IV, the proposed algorithm demonstrates significant advantages in both operational cost and voltage violation rate. Specifically, the average cost reached 4840.61 CNY, while the voltage violation rate was $1.06\times 10^{-3}$. Through comprehensive analysis of mean-standard deviation comparisons with baseline algorithms, the proposed approach achieves optimal results across all three evaluation metrics. It maintains the lowest mean values while exhibiting minimal standard deviations, indicating not only superior average performance but also robust stability.
@@
During the training process, the evolution of the agent average Lagrange multiplier $\lambda$ and the average Wasserstein-2 metric provides key insight into the dynamic balance between expert imitation and autonomous policy optimization. As shown in Fig.7, $\lambda$ initially increases rapidly, reflecting the strong constraint effect imposed by the LLM expert strategy to guide the early-stage exploration of the MARL agents. With the gradual improvement of policy learning, $\lambda$ subsequently decreases and eventually stabilizes, indicating that the agents have progressively internalized the expert behavior and require less external constraint. In contrast, the Wasserstein-2 metric between the agent policy and the expert policy exhibits a consistently decreasing trend, demonstrating continuous convergence of the learned policy toward the expert distribution. Once the Wasserstein-2 metric approaches the predefined threshold Policy deviation $\epsilon$, it remains stable, signifying that the imitation learning constraint has reached equilibrium and the training has achieved convergence.
@@
To further investigate the sensitivity of the Wasserstein-based regularization, a comparative analysis was conducted between two formulations of the Wasserstein-2 metric in (34): one with the variance term $\sigma$ and one without it. As illustrated in Fig.8, when the $\sigma$ is incorporated, it converges smoothly to approximately $10^{-5}$ after around 500 episodes. This indicates that the stochastic exploration of the actor network is gradually reduced as the learned policy approaches the expert strategy, achieving a stable and deterministic behavior. In contrast, when the $\sigma$ is excluded, the implied policy variance remains large and fluctuates around 1.2 throughout training, suggesting incomplete convergence, the agent’s actions are not stable enough. These results highlight that explicitly learning and regularizing the $\sigma$ is crucial for maintaining exploration in the early stage while ensuring eventual convergence to a stable policy distribution.
@@
Figure 9: Ablation study on the number of attention heads
To evaluate the effectiveness of the proposed differential attention mechanism in multi-agent P2P energy trading, ablation studies were conducted, with results shown in Fig. 9. Experimental findings reveal that the number of attention heads significantly impacts cost deviation in LLM expert. When the number of attention heads increases from a small value (e.g., from 0 to 1 or from 1 to 2), the system exhibits a significant reduction in cost deviation. This indicates that the introduction of a few attention heads provides substantial gains in capturing key agent interactions. However, as the number of heads increases to a higher level (e.g., from 8 to 16), the performance improvement tends to saturate, and the reduction in cost deviation becomes marginal. This suggests that an excessive number of attention heads may introduce redundant information, thereby suppressing model performance. Furthermore, under the configuration where the head count satisfies $d_{model}/2d$, the differential attention mechanism demonstrates superior scheduling performance while maintaining equivalent parameter scale and computational complexity compared to standard attention mechanisms.
It is worth emphasizing that the precise optimization models generated by the LLM are only employed during the training phase, assuming access to future information, and therefore cannot be directly applied in real-world real-time P2P energy market environments. In our framework, the LLM functions solely as an expert sample generator in the training stage, which accelerates the convergence of MARL and mitigates the issues of slow convergence and high exploration cost inherent in reinforcement learning. After training is completed, the MARL policy network can make rapid decisions relying solely on the current state, thereby fully meeting the stringent timeliness requirements of real-time P2P energy markets.

### v3

@@ III-B1 Preparation
The expert strategy constraints are embedded within a multi-agent actor-critic framework to solve the constrained optimization problem. Based on the Lagrangian dual theory, this problem is first transformed into its Lagrangian dual form. For each prosumer agent $i$, the multi-agent formulation can be expressed as:
@@
where $\mathcal{B}$ is the experience replay buffer; $\phi$ is the policy network parameters and the output sampling from the Gaussian distribution $a_{i,t}\sim\mathcal{N}(\mu_{\phi_{i}},\sigma_{\phi_{i}}^{2})$; $\epsilon$ is the policy deviation; $\theta$ is the Q-function network parameters and z = 1,2; $\psi$ is the V-function network parameters. The policy network outputs a distribution over actions, which inherently facilitates exploration during training by enabling diverse action sampling under identical states.
@@
The resulting Wasserstein-2 metric between the expert strategy and the policy network is:
@@ III-B2 Multi-Head Differential Attention
To enhance the modeling capability of the centralized critic in capturing inter-agent interactions, a multi-head differential attention mechanism is integrated into the critic network. This design aims to improve the accuracy of global value estimation. For each agent $i$, the input is first transformed into an embedding $e_{i}$ and then $E$ represents the traversal of these embedded vectors. Matrix $E$ is then split into $h$ attention heads and linearly projected into the query, key, and value spaces for each head, as follows:
@@
To enhance the representational capacity of critic networks while preserving ego-specific information, we employ a feature fusion mechanism that concatenates the intermediate output $x_{i}$ with the original agent embedding $e_{i}$. This ensures that agent-specific features remain directly accessible to subsequent layers, mitigating potential information loss during intermediate transformations. The fused representation is processed by an MLP of fixed dimensionality, guaranteeing that its size remains invariant to agent population growth. This fusion facilitates flexible feature interactions and refines the representation prior to generating the critic’s final value estimate. The overall architecture is illustrated in Fig. 3.
@@ III-B3 Learning the Critics
In scenarios where agents share a global reward, a major challenge is to reduce the variance of policy gradient estimates in interactive multi-agent environments. To address this issue, a double Q-function network and a target V-function network are employed. The minimum selection operation in the double Q-function network effectively mitigates overestimation bias. The incorporation of a V-function network contributes to a significant reduction in estimation variance [47]. All agents share critic networks with shared parameters; this design significantly reduces the number of trainable parameters and improves training stability and scalability in large-scale multi-agent systems.
For V-function network $V_{\psi}$, the loss function is calculated to approximate the given the current state and Q-function values:
$\mathcal{L}_{V}(\psi)=\mathbb{E}_{s_{t}\sim\mathcal{B},{a}_{t}\sim\pi_{\phi_{i}}(\cdot|{{o}_{i,t}})}\Bigl[(V_{\psi}({s}_{t})-\underset{z\in{1,2}}{\text{min}}Q_{\theta_{z}}(s_{t},{a}_{t}))^{2}\Bigr].$ (39)
For the Q-function network $Q_{\theta}$, the training target $y_{t}$ is computed using the a delay updated target network $V_{{\overline{\psi}}}$. The loss function is defined as follows:
$\begin{split}\mathcal{L}_{Q}(\theta)=&\mathbb{E}_{(s_{t},\mathbf{a}_{t},r_{t},s_{t+1})\sim\mathcal{B}}\biggl[\frac{1}{2}\left(Q_{\theta}(s_{t},a_{t})\right.\left.-y_{t}\right)^{2}\biggr]\\ &y_{t}=r_{i,t}+\gamma V_{\bar{\psi}}(s_{t+1})\ ,\end{split}$ (40)
where the target V-function network $\overline{\psi}$}} is updated via Polyak averaging rather than direct copying to enhance stability:
$\bar{\psi}\leftarrow\tau\psi+(1-\tau)\bar{\psi}\ ,$ (41)
@@ III-B4 Learning the Actors
In the initial phase of training, a low policy deviation is employed to guide the agent’s learning; during the middle and later phases, a larger policy deviation is introduced to sustain the agent’s exploration. This framework ultimately enables efficient and stable policy optimization.
@@ IV NUMERICAL STUDY
To validate the effectiveness of the proposed framework, a numerical study is carried out on a modified IEEE 141‑bus distribution network, the voltage level is 12.47kV. As shown in Fig. 4, 20 prosumers are selected to participate in local P2P energy trading, and the characteristics of five prosumer types are summarized in TABLE II. The renewable generation outputs and load demand curves of the prosumers are extracted from a real‑world dataset published by the Belgian Transmission System Operator Elia[49].
Node Devices Portfolio Prosumer Scenario CDG WT PV BESS CL 48,78,102,127, Commercial 59,109,130,140 Rural 67,95,133,136 Industrial 62,86,106,138 Residential 74,100,116,134 Energy Hub
TABLE II: Personalized configurations for 20 prosumers
@@ IV-A Comparison Baselines
    MAGAIL[53]: A multi-agent extension of GAIL, in which an adversarial discriminator is trained to distinguish expert trajectories generated by LLM from the agents’ joint state–action trajectories. The output of the discriminator is subsequently transformed into a shaped imitation reward, which replaces the environment reward.
@@
    Our Proposed: Introduces a Lagrange multiplier during training to progressively constrain agent behaviors toward the LLM expert’s strategies.
@@ IV-B Implementation Details
In terms of neural network architecture design, aside from the variants incorporating attention mechanisms, all baseline algorithms share a unified network architecture under identical hyperparameter settings. The detailed hyperparameter settings are provided in TABLE III. Training was conducted for 5000 episodes with the same random seed initialization and update frequency. All experiments were implemented in Python 3.11.10 under the PyTorch 2.7, with parameters updated via the Adam optimizer. Computations were performed on a platform equipped with an NVIDIA RTX 5070Ti GPU, an AMD Ryzen ThreadRipper 3970X CPU, and 64GB of RAM.
TABLE III: Hyperparameters used in the MARL
@@
Using year-round data, we designate the first day of each month as the validation set, the last week of each month as the test set, and the remaining days for training. A dynamic validation strategy is employed during training. After each training step, the current policy is evaluated using the validation set. The average cumulative discounted reward across all validation sets is recorded as a performance metric. To ensure statistical robustness, each algorithm is executed independently five times, and both the mean reward and the standard deviation of the resulting rewards are recorded.
In the present study, the large parameter LLM workflow is implemented through online API interfaces for experimental validation. Nevertheless, the proposed architecture is fully modular, allowing seamless substitution with locally hosted or SFT LLM instances to ensure data privacy and reduce latency. The LLM workflow is activated after every 200 episodes during the MARL training phases or system reconfiguration stages (e.g., when new devices are added or topology changes occur). This approach effectively mitigates the decline in learning performance caused by any poor generation by the LLM in a given episode. During every training step, each prosumer executes the LLM pre-generated optimization model code in parallel, while the DSO merely aggregates the objective functions and constraints from all prosumers and utilizes a commercial solver to generate secure operational strategies for each prosumer. To address potential runtime burdens, efficient memory management is implemented to ensure the computational tractability of the iterative training process. In real-time operation, lightweight MARL agents autonomously perform decision-making without additional LLM guidance or global DSO validation. In particular, when the LLM expert workflow encounters a device that is not available in the knowledge base, the policy deviation is deliberately increased to a very large value. Under this setting, the agent in our proposed algorithm that lacks valid LLM expert demonstrations effectively degenerates into a reward-driven reinforcement learning process, as the influence of expert guidance is implicitly suppressed.
@@ IV-C1 Analysis of the LLM Expert Strategy
In this evaluation, the human expert baseline refers to the optimal strategy obtained by solving a deterministic convex optimization model—formulated in accordance with the problem setting described in Section II—using the Gurobi. This model fully incorporates all physical constraints of the prosumers and minimizes the total operational cost under perfect forecasts of renewable generation and load demand. The resulting solver-based optimal policy serves as the ground-truth reference for assessing the accuracy of LLM-generated strategies.
@@
TABLE IV: Performance comparison of different LLMs in workflow without prosumer preference
@@
TABLE IV provides a detailed comparison of the four metrics proposed in this study, where the LLMs above the dashed line deactivates advanced reasoning capabilities, while those below activate it. The latency of a large model depends on the network speed when API calls are used, but is effectively eliminated when the model is deployed locally. Specifically, LLMs without advanced reasoning capabilities are generally more cost-efficient; among these, Google Gemini-2.5-Flash shows the highest average cost at 2.43 CNY (Derived from the official billing statements of different LLMs after running the workflow once, and converted into CNY based on the prevailing exchange rate), mainly because it generates a larger number of output tokens per query, leading to higher overall billing despite its relatively low per-token price. In contrast, among reasoning-enabled models, Claude 4-Opus is the most expensive, with an average price of 6.27 CNY, primarily due to its higher per-token pricing structure reflecting the premium positioning of its advanced reasoning capabilities. Open-source LLMs such as DeepSeek and Qwen can also completely eliminate API-related expenses when deployed locally.
@@
Figure 5: Comparative experiments on the LLM framework without prosumer preference
@@
To further investigate the performance of personalized prosumer preferences within the LLM expert workflow, we consider the potential conflicts between prosumer preferences and distribution network voltage constraints. Accordingly, the prompt instructs the LLM to formulate prosumer preferences as soft constraints in the objective function. This design ensures that distribution network security is always prioritized when conflicts arise, while prosumer preferences are satisfied to the greatest extent possible. The results are illustrated in Fig. 6.
Figure 6: Our proposed LLM expert workflows for processing personalized prosumer preferences
TABLE V: Performance comparison of different LLMs in workflow with prosumer preference
LLM Pass Rate(%) Accuracy(%) Correction Tokens Chatgpt-4o 86 92.27 1.56 4904 Claude-3.5-Sonnet 92 99.14 1.01 6223 Gemini-2.5-Flash 90 98.34 1.47 19341 DeepSeek-V3 $\bm{96}$ 96.01 $\bm{0.54}$ $\bm{4362}$ Qwen-2.5-Max 88 $\bm{99.39}$ 1.75 5638 Chatgpt-o3 $\bm{100}$ 98.35 0.38 17773 Claude-4-Opus $\bm{100}$ $\bm{99.81}$ $\bm{0.28}$ 13320 Gemini-2.5-pro $\bm{100}$ 99.79 0.57 $\bm{8365}$ DeepSeek-R1 $\bm{100}$ 98.01 0.65 17322 Qwen-3 $\bm{100}$ 99.66 0.43 9010
TABLE V evaluates the performance of LLMs in handling prosumer preferences using 50 randomly generated, unambiguous profiles per prosumer type to ensure a fair comparison with human experts. Compared to TABLE IV, the Pass Rate remains unchanged, while the Accuracy decreases slightly. This slight decline occurs because incorporating preferences increases the workflow complexity, necessitating more iterative refinements; consequently, both the Correction metric and Token count increase, although this does not compromise the LLMs’ ability to successfully generate executable code. Further demonstrating the method’s robustness, the results in Fig. 7 highlight the advantages of the proposed workflow when handling the increased complexity of prosumer-defined optimization models. In contrast to Fig. 5, omitting the proposed workflow in this setting leads to significant degradation: the lowest Pass Rate drops to only 30%, while Accuracy falls to 85.43%, proving the workflow’s superior effectiveness in managing tasks with heightened complexity.
Figure 7: Comparative experiments on the LLM framework with prosumer preference
TABLE VI: Performance Comparison of LLMs across Different Parameter in workflow with prosumer preference
LLM Pass Rate(%) Accuracy(%) Correction Tokens without preferences Qwen-3-32B 82 89.34 2.21 6361 Qwen-3-14B 76 85.63 4.03 6908 Qwen-3-8B 44 76.85 1.27 6005 Qwen-3-4B 2 82.47 1.78 6187 with preferences Qwen-3-32B 78 87.41 2.63 6398 Qwen-3-14B 64 85.08 4.17 7036 Qwen-3-8B 26 74.86 1.61 6121 Qwen-3-4B 0 78.12 1.74 6202
To address API connection limitations, we evaluated locally deployed open-source Qwen-3 models with reasoning activated across various parameter scales in TABLE VI, testing 50 tasks with user preferences and 50 without. Workflow effectiveness declines with model size; reducing the scale from 14B to 8B makes matching human-expert Pass Rate and Accuracy unattainable.LangGraph workflow analysis reveals that models under 8B parameters consistently fail to transform input specifications. These models omit key user devices, preferences, and Model Generation Expert-derived constraints, while also failing to select appropriate RAG-retrieved Atomic Functions. Furthermore, these models have nearly lost the capability to rectify erroneous code. Consequently, in successful instances, the number of correction rounds and token consumption actually decrease due to the lack of iterative refinement. Thus, Qwen-3 models with at least 14B parameters are recommended for local deployment.
@@ IV-C2 Performance Analysis of MARL Algorithm
During the training of 20 prosumers, the proposed LLM expert workflow average replaced 3,039 lines of manually written human expert code. The results in Fig. 8 demonstrate the performance of the proposed algorithm during training on the validation set. As shown, ’Our proposed-MH’ achieves faster convergence to a low operational cost with minimal fluctuations, demonstrating enhanced reward stability for guiding agents when LLM outputs expert actions. Furthermore, the curve of average voltage violation rate for ’Our proposed-MH’ rapidly declines and maintains a low level, highlighting its strength in constraint satisfaction and ability to maintain secure grid operations.
TABLE VII: Comparison of different algorithms on test set
@@
After completing model training, we applied the proposed algorithm to a test set to evaluate the practical applicability. As shown in TABLE VII, the proposed algorithm demonstrates significant advantages in both operational cost and voltage violation rate. Specifically, the average cost reached 4840.61 CNY, while the voltage violation rate was $1.06\times 10^{-3}$. Through comprehensive analysis of mean-standard deviation comparisons with baseline algorithms, the proposed approach achieves optimal results across all three evaluation metrics. It maintains the lowest mean values while exhibiting minimal standard deviations, indicating not only superior average performance but also robust stability.
@@
During the training process, the evolution of the agent average Lagrange multiplier $\lambda$ and the average Wasserstein-2 metric provides key insight into the dynamic balance between expert imitation and autonomous policy optimization. As shown in Fig. 9, $\lambda$ initially increases rapidly, reflecting the strong constraint effect imposed by the LLM expert strategy to guide the early-phase exploration of the MARL agents. With the gradual improvement of policy learning, $\lambda$ subsequently decreases and eventually stabilizes, indicating that the agents have sufficiently aligned with the expert behavior and require less external constraint. In contrast, the Wasserstein-2 metric between the agent policy and the expert policy exhibits a consistently decreasing trend, demonstrating continuous convergence of the learned policy toward the expert distribution. Once the Wasserstein-2 metric approaches the predefined threshold Policy deviation $\epsilon$, fluctuations remain bounded, signifying that the imitation constraint has reached equilibrium and the policy optimization process has entered a stable regime.
@@
To further investigate the sensitivity of the Wasserstein-based regularization, a comparative analysis was conducted between two formulations of the Wasserstein-2 metric in (34): one with the variance term $\sigma$ and one without it. As illustrated in Fig. 10, when the $\sigma$ is incorporated, it converges smoothly to approximately $10^{-5}$ after around 500 episodes. This indicates that the stochastic exploration of the policy network is gradually reduced as the learned policy approaches the expert strategy, achieving a stable and near-deterministic policy behavior. In contrast, when the $\sigma$ is excluded, the implied policy variance remains large and fluctuates around 1.2 throughout training, suggesting incomplete convergence, the agent’s actions are not stable enough. These results highlight that explicitly learning and regularizing the $\sigma$ is crucial for maintaining exploration in the early phase while ensuring eventual convergence to a stable policy distribution.
To investigate the prosumer preference issue in P2P energy trading, heterogeneous preferences are assigned to 20 prosumers on the IEEE 141-bus topology shown in Fig. 4. For each prosumer’s personalized preference, Our proposed method is guided by an LLM-based expert workflow.
Figure 11: The operate results of our proposed algorithm on BESS and CL preference prosumer
The results in Fig. 11 present a comparative analysis for prosumers exhibiting preferences toward BESS and CL operation under the proposed method and MAAC. Conventional MARL algorithms, represented by MAAC, are limited to conventional optimization behavior due to their strict reliance on the predefined reward function. In contrast, the proposed method further aligns the learned policy with expert strategies beyond purely reward-driven optimization, thereby enabling the consideration of prosumer preferences while simultaneously maintaining the maximization of social welfare.
When the device type requested by a prosumer is not included in the knowledge base, or in local deployment settings, the use of LLMs with fewer than 8B parameters may lead to suboptimal expert actions. A comparative experiment is conducted between the proposed method and the existing baseline algorithm under a Qwen-3-8B model. After 1,000 training episodes, we explicitly schedule an increase in the policy deviation to allow the algorithm to perform autonomous exploration, while enforcing a lower bound policy variance to prevent premature policy determinism.
Figure 12: Training results of baseline algorithm under Qwen-3-8B strategies
As shown in Fig. 12, conventional imitation learning (e.g., BC, GAIL) excessively constrains policies, causing agents to inherit expert bias and converge to inferior local optimal. In contrast, our method gradually relaxes the imitation constraint by increasing policy deviation. This expands the exploration space and reintroduces reward signals, enabling agents to escape LLM expert-induced suboptimal equilibrium. The early-phase alignment provides stable initialization, while later-phase self-exploration compensates for expert inaccuracies. This two-phase process effectively balances expert knowledge exploitation with autonomous exploration. Conversely, BC- and GAIL-based methods enforce strict adherence throughout, rendering them unable to recover from suboptimal expert guidance.
To assess the scalability of Our proposed-MH algorithm under large-scale prosumer participation, further experiments are conducted on the IEEE 141-bus distribution network with 100 prosumers, the prosumer details are reported in TABLE VIII.
TABLE VIII: Personalized configurations for 100 prosumers on IEEE-141 system
Node CDG WT PV BESS CL Prosumer Type 10, 12, 18, 24, 28, 33 Commercial 45, 48, 61, 78, 94, 102 15, 21, 30, 37, 52, 59 Rural 71, 84, 96, 109, 123, 130 17, 26, 36, 44, 54, 67 Industrial 75, 88, 95, 111, 124, 133 19, 29, 39, 47, 57, 62 Residential 76, 86, 99, 106, 121, 138 22, 31, 41, 50, 60, 74 Energy Hub 83, 100, 112, 116, 128, 134 6, 13, 23, 34, 46, 58, 68 Renewable 82, 97, 108, 114, 126, 135, 141 9, 16, 27, 35, 49, 57, 70 Storage-centric 81, 91, 101, 113, 125, 139 38, 56, 63, 72, 80, 90, 103 Data Center 110, 118, 129, 131, 136, 140
Figure 13: Comparison chart of the performance of baseline algorithms with large-scale prosumers
Fig. 13 demonstrates the proposed method’s scalability, with the LLM workflow replacing 14,874 lines of manual code. Conventional CTDE baselines (e.g., MATD3+BC, MAGAIL) fail in this setting as they require training 100 distinct critics with linearly growing input dimensions. In contrast, our proposed-MH algorithm addresses these limitations using a shared critic and Differential Attention, which ensures the concatenated representation ($Concat(e_{i},x_{i})$) remains fixed regardless of system scale. Together with agent-specific LLM guidance, this guarantees effectiveness in large-scale scenarios.
@@
Figure 14: Ablation study on the number of attention heads in 20 agent prosumer
To evaluate the effectiveness of the proposed differential attention mechanism in multi-agent P2P energy trading, ablation studies were conducted, with results shown in Fig. 14. The experimental results indicate that the number of attention heads substantially affects the cost deviation in the LLM expert. Increasing the number of heads from very few (e.g., 1–2) leads to a notable reduction in cost deviation, suggesting that introducing a small number of attention heads significantly enhances the model’s ability to capture key agent interactions. However, further increasing the head count (e.g., from 8 to 16) yields only marginal improvements, indicating performance saturation likely due to redundancy in the captured features. Crucially, the differential attention mechanism outperforms standard attention in scheduling accuracy while preserving similar computational complexity. Its two sequential softmax operations incur roughly the same cost as a standard multi-head attention with twice the number of heads, thus delivering superior performance without computational overhead.
TABLE IX: Ablation Study of Initial $\xi$
Initial $\xi$ Episode 100 Episode 1000 Episode 5000
Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
0.2 -82.65 3.74 -20.54 3.27 -16.56 1.17
0.5 -85.23 4.13 -20.20 3.31 -17.30 1.22
0.8 -92.91 3.95 -21.16 2.99 -16.89 1.24
An ablation study is conducted to investigate the impact of the initialization of the learnable vector $\xi$ in the Differential Attention mechanism. As shown in TABLE IX, the choice of the initial value has only a minor effect on training performance during the early phase (within the first 100 training steps). After 1,000 training steps, the initialization of $\xi$ has a negligible influence on both the average training reward and its variance. These results indicate that $\xi$ can consistently adapt during training and gradually overcome the bias introduced by its initial value, demonstrating the robustness of the proposed Differential Attention mechanism with respect to parameter initialization.
It is important to note that existing RL methods typically separate the training phase from real-world deployment, utilizing simulation environments rather than interacting directly with the real-world power system[12]. In this simulation-based training phase, traditional optimization methods are feasible as global information is accessible. However, they are often infeasible for real-time deployment due to the lack of accurate forecast parameters [16]. Consequently, the LLM-generated optimization models are employed strictly to accelerate MARL training convergence. The trained agents are capable of autonomously executing energy trading tasks based solely on local observations, adaptively handling real-time uncertainties without further LLM intervention.